# MERLIN: Multi-stagE query performance prediction for dynamic paRallel oLap pIpeliNe

### v1

@@ Abstract.
High-performance OLAP database technology has emerged with the growing demand for massive data analysis. To achieve much higher performance, many DBMSs adopt sophisticated designs including SIMD operators, parallel execution, and dynamic pipeline modification. However, such advanced OLAP query execution mechanisms still lack targeted Query Performance Prediction (QPP) methods because most existing methods target conventional tree-shaped query plans and static serial executors. To address this problem, in this paper, we proposed MERLIN a multi-stage query performance prediction method for high-performance OLAP DBMSs. MERLIN first establishes resource cost models for each physical operator. Then, it constructs a DAG that consists of a data-flow tree backbone and resource competition relationships among concurrent operators. After using a GAT with an extra attention mechanism to calibrate the cost, the cost vector tree is extracted and summarized by a TCN, ultimately enabling effective query performance prediction. Experimental results demonstrate that MERLIN yields higher performance prediction precision than existing methods.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
1. Introduction
@@
OnLine Analytical Processing (OLAP) databases are a type of database system designed specifically to support complex queries and large-scale data analysis. With the development of OLAP DBMSs, there were already many advanced execution mechanisms proposed to improve OLAP DBMS’s execution speed, such as applying the bulk (chunk) processing to speed up queries used by PostgreSQL and MonetDB (Boncz et al., 2005). Some advancing OLAP DBMSs such as ClickHouse (Schulze et al., [n.d.]), SparkSQL (Lyu et al., 2024), Databricks (Xue et al., 2024), and DuckDB (Raasveldt and Mühleisen, 2019) pushed the OLAP query processing technology to a new level by introducing a dynamic pipeline instead of using a conventional static query execution engine. Some of them such as ClickHouse and SparkSQL go a step further and use a highly parallel DAG-based pipeline structure instead of the tree-shape physical plan executed by a Volcano-style engine. Such huge and unique modifications to the query plan model and engine lead to many new demands for components like the cost estimator and query optimizer etc. As shown in Table 1, we refer to such OLAP DBMSs support all the mechanisms as high-performance OLAP DBMSs and focus on the learned QPP method toward them.
@@
Table 1. Comparison of mechanisms supported by DBMSs. S: support, N: not-support, PS: partial-support, IS: indirect-support, AQE: adaptive query execution, and Dynamic: dynamically changing operators and pipelines. Exe. is the abbreviation for Execution.
@@
Query performance prediction is a widely needed task in admission control (Tozer et al., 2010; Wu et al., 2013b), resource management (Wu et al., 2013b; Chi et al., 2011; Taft et al., 2016), and query optimization (Marcus et al., 2019; Lan et al., 2021; Lyu et al., 2024; Xue et al., 2024), which mainly aims to estimate the latency of queries. The accuracy of QPP directly impacts all the above data management tasks and is crucial to the performance and stability of DBMS. QPP can be regarded as a specific type of cost estimation problem that uses query latency to measure query cost. Before the widespread introduction of deep learning into the database domain, mainstream databases primarily employed conventional methods to predict query plan performance. These methods mostly relied on simple linear models, combining the cost coefficients of physical operators such as sequential scan and indexing with the input scale provided by cardinality estimators to derive query performance. Although such conventional methods are widely used in some advanced OLAP DBMSs such as SparkSQL (Lyu et al., 2024) and Databricks (Xue et al., 2024), they ignore the inherent complexity of DBMS and can only achieve limited accuracy. For example, the single operator’s cost can be variable under different query pipeline contexts and different system workloads due to the cache changing and resource utilization caused by other operators, especially under the dynamic pipeline environment. Such complexity is even more significant in OLAP DBMSs such as ClickHouse, which takes the high degree of parallelism of query pipelines as the core design idea. Solving such complex problems is exactly what the learning cost estimator is good at.
@@
With the development of deep learning, numerous studies have applied deep learning models to the QPP field, aiming to directly measure the quality of query plans based on the time cost of query execution. Existing works primarily focus on feature representation and end-to-end prediction for tree-shaped query plans. For instance, in databases like PostgreSQL (Stonebraker and Rowe, 1986) where query plan trees serve as physical plans, many methods (Marcus et al., 2019; Sun and Li, 2019; Marcus and Papaemmanouil, 2019) have been demonstrated to achieve higher accuracy and prediction speed. In the field of concurrent query performance prediction, there have also been study (Zhou et al., 2020) treating concurrent queries’ plan trees as Directed Acyclic Graphs (DAGs) and treating resource competition and data sharing relationships between concurrent queries as edges that connect those sub-DAGs to capture resource competition relationships among a group of operators.
However, most existing learned methods are designed for DBMSs with tree-like query plans and static serial pipelines. For high-performance OLAP DBMSs that use SIMD operators, DAG-based pipelines, and dynamic parallel execution engines such as ClickHouse, DuckDB, and SparkSQL, there is still a lack of targeted learned QPP methods. Therefore, it is urgent to study the corresponding learning cost estimation method for such OLAP DBMSs with dynamic parallel pipelines. We summarized three main challenges for such DBMSs’ query performance prediction:
C.1 SIMD Operators’ resource prediction. More and more DBMSs (Schulze et al., [n.d.]; Kornacker et al., 2015; Boncz et al., 2005) use SIMD instructions to implement physical operators, whose cost variance is larger under different input scales and system resource utilization rates, and their resource costs are more difficult to predict but also important to the overall query’s performance prediction. Common methods of predicting query performance end-to-end used by most existing approaches cannot capture each operator’s cost effectively enough. A new process needs to be proposed to make a more fine-grained estimate of operator resource overhead.
@@
C.2 Dynamic pipeline extraction and expression. Some DBMSs (Lyu et al., 2024; Xue et al., 2024; Schulze et al., [n.d.]) employ runtime optimization to alleviate issues such as inaccurate estimation of key information such as cardinality before querying, or changes in data distribution and system load during runtime. Among all these DBMSs, ClickHouse uses the most unique mechanism that allows dynamic modifications on the pipeline itself. For example, ClickHouse will add extra operators during execution, and change the algorithm of operators such as sort and join. In summary, there are significant differences between ClickHouse and other databases, and existing QPP methods are not designed for such a unique framework. Such a mechanism makes it inaccurate to use the pipeline extracted before the query, and it is also difficult to express any dynamic pipeline extracted to a form that can be easily processed by learned QPP methods. It’s difficult to extract a low-level runtime dynamic pipeline that is detailed enough to support the QPP task. Also, even if such information can be extracted, it is still difficult to process with low cost by learned QPP methods, especially when the pipeline is in spatiotemporal graph form since the inference on spatiotemporal graph networks is expensive.
@@
C.3 Parallel execution and resource competition. DBMSs such as ClickHouse and SparkSQL employ parallel pipeline executors, which introduces more complex resource competition within a single OLAP query. Taking ClickHouse  [^1] https://github.com/ClickHouse/ClickHouse/issues/34045 as an example, its parallel execution of ClickHouse consists of two parts, horizontal and vertical parallel. The horizontal parallel execution means that different parts of data are processed in parallel by the same computation. ClickHouse achieves this by widely using SIMD instruction sets like SSE and AVX  [^2] https://clickhouse.com/blog/cpu-dispatch-in-clickhouse and creating multiple execution branches (threads) for operators like MergeSorting to speed up queries. Such operators’ costs are highly related to the system workload and much harder to predict within an end-to-end model. The vertical parallel execution means that different stages of computation are run in parallel. To fit such a design and speed up queries, different from the widely used volcano model, ClickHouse’s query plans are constructed as parallel pipelines [^3] https://presentations.clickhouse.com/meetup24/5.%20Clickhouse%20query%20execution%20pipeline%20changes/, resulting in significant resource competition among operators within a single query plan, including competition for CPU, I/O, and memory resources. Existing methods focus on common databases with tree-like query plans, how to capture those resource competition information and calibrate operators’ costs still remains to study.
@@
In this paper, we propose a multi-stage QPP method MERLIN to tackle the challenges above and give an accurate and efficient prediction, which decouples the operator cost prediction, cost calibration, and query performance prediction. The core idea of MERLIN to achieve high accuracy for high-performance OLAP DBMSs is first to obtain detailed low-level physical pipeline information that includes dynamic pipeline changes and then to use multi-stage models to predict the resource cost of SIMD operators separately, calibrate the resource competition, and summarize the overall pipeline latency.
@@
For challenge C.1, MERLIN decouples operator cost prediction from query performance prediction. We establish lightweight cost models (a.k.a operator cost predictors (OCPs)) for each operator to predict the resource and time cost of each operator’s execution as Stage 1. These individual operator cost vectors are then aggregated into a graph for subsequent processing.
@@
For challenge C.2, MERLIN includes a Runtime Tracker to extract training data and necessary dynamic information for prediction. To collect training data of the OCPs, MERLIN uses a novel serial executor and the tracker runs on the full-collection mode that marks all data chunks by their addresses. For query-level training data of training and prediction, considering high-performance OLAP DBMSs’ dynamic parallel execution mechanism and inspired by ClickHouse’s probe phase (Schulze et al., [n.d.]), the tracker uses a probe execution mode which only launches a few chunks to obtain operator features and the execution paths, which usually contain more operators than the initial pipeline. For dynamic pipeline expression, MERLIN uses the data-flow tree rather than the pipeline DAG as the backbone of the data structure. MERLIN includes a Graph Constructor to build the resource competition graph with the backbone data-flow tree and corresponding operators’ cost vectors based on the execution paths of chunks collected by the Runtime Tracker. Therefore, the chunks’ execution information generated by the modified pipeline can be reserved in different execution paths of the data-flow tree without mixing up.
@@
For challenge C.3, we introduced resource competition into the pipeline and designed an attention-based resource competition intensity adaptive mechanism and then used Graph Attention Network(GAT) (Veličković et al., 2017) to calibrate the cost as Stage 2. In particular, MERLIN categorizes operators based on their resource usage and assigns them to different resource competition groups. We establish edges in the generated data-flow tree based on those resource competition groups to represent resource competition relationships among operators within the same group. The edges’ weights are self-adjusted by an attention mechanism that is coupled with the data flow structure. Next, we use a GAT model to calibrate operator costs. Finally, to gather those calibrated costs into the final QPP result, we develop a method to convert the graph back to a tree in a differentiable way and use a TCN to aggregate those costs from the bottom to the top and predict the query performance as Stage 3.
@@
    We design a multi-stage QPP approach that decouples the operator cost estimation, resource competition calibration, and query performance prediction to provide a more accurate QPP for high-performance OLAP DBMSs. It fills a gap in the field of learned cost estimators of high-performance OLAP DBMSs with dynamic parallel DAG pipelines.
  2. (2)
@@
    We design a low-level Runtime Tracker and implement it on ClickHouse to collect detailed and accurate data for the QPP model’s training and prediction, which includes full-collection mode, probe execution mode, Cost Logger and a modified serial executor.
  3. (3)
@@
    Experimental results demonstrate that MERLIN achieves higher estimation accuracy and comparable or lower time overhead than baseline methods.
2. Related Works
2.1. Conventional Methods
@@
The most common method is to build cost models for different operators before learned methods are introduced into the database system. Such methods employ the CPU cost, I/O cost, etc. of operators to estimate the query cost. Manegold et al. proposed a series of cost functions to different memory access patterns (Manegold et al., 2002). These functions are parameterized to fit different hardware features. The operators’ cost can be inferred from the combination of those cost functions. Feilong Liu et al. proposed a memory I/O cost model (Liu and Blanas, 2015) to identify good evaluation strategies for complex query plans with multiple hash-based equip-joins over memory-resident data. Wentao et al. proposed a cost model (Wu et al., 2014) that represents the query latency as a function of the selectivities of operators in the query plan as well as the constants that describe the cost of CPU and I/O. They solve the coefficient of the cost model from the observed query information. Wu et al. (Wu et al., 2013a) proposed an offline profiling method to set the coefficients of a cost model for specific hardware and software conditions.
@@
Such methods have two main problems, the first problem is that the parameters of the cost models are hard to adjust properly to maximize accuracy. Although there are some methods  (Wu et al., 2014; Vasilenko et al., 2024) proposed to adjust or solve the parameters, it’s not accurate enough for dynamic workload environments since the ground truth costs change with the hardware workload. The second problem is that some research  (Leis et al., 2015; Sun and Li, 2019) show that the cardinality estimation results are crucial to those cost models, which is another harder problem that has much higher error usually. One important reason is that compared to learned models, conventional methods’ prediction results highly depend on the input scale and lack of robustness like learned methods.
2.2. Learned Methods

### v2

# CONCERTO: Complex Query Execution Mechanism-Aware Learned Cost Estimation

@@ Abstract
With the growing demand for massive data analysis, many DBMSs have adopted complex underlying query execution mechanisms, including vectorized operators, parallel execution, and dynamic pipeline modifications. However, there remains a lack of targeted Query Performance Prediction (QPP) methods for these complex execution mechanisms and their interactions, as most existing approaches focus on traditional tree-shaped query plans and static serial executors. To address this challenge, this paper proposes CONCERTO, a Complex query executiON meChanism-awaE leaRned cosT estimatiOn method. CONCERTO first establishes independent resource cost models for each physical operator. It then constructs a Directed Acyclic Graph (DAG) consisting of a dataflow tree backbone and resource competition relationships among concurrent operators. After calibrating the cost impact of parallel operator execution using Graph Attention Networks (GATs) with additional attention mechanisms, CONCERTO extracts and aggregates cost vector trees through Temporal Convolutional Networks (TCNs), ultimately achieving effective query performance prediction. Experimental results demonstrate that CONCERTO achieves higher prediction accuracy than existing methods.
Index Terms:
Database Management, AI4DB, Cost Estimation, Query Performance Prediction
I Introduction
@@
Online Analytical Processing (OLAP) databases are a type of database system designed specifically to support complex queries and large-scale data analysis. With the development of OLAP DBMSs, there were already many advanced execution mechanisms proposed to improve OLAP DBMSs’ execution speed, such as applying the bulk (chunk) processing to speed up queries used by PostgreSQL and MonetDB [31]. Some advanced OLAP DBMSs such as ClickHouse [20], SparkSQL [39], Databricks [38], and DuckDB [32] pushed the OLAP query processing technology to a new level by introducing a dynamic pipeline instead of using a conventional static query execution engine. Some of them, such as ClickHouse and SparkSQL, go a step further and use a highly parallel DAG-based pipeline structure instead of the tree-shaped physical plan executed by a Volcano-style engine. Such huge and unique modifications to the query plan model and engine lead to many new demands for components like the cost estimator and query optimizer, etc. As shown in Table I, we refer to OLAP DBMSs that support all these complex execution mechanisms as high-performance OLAP DBMSs and focus on the learned QPP method toward them.
@@
TABLE I: Comparison of mechanisms supported by DBMSs. S: support, N: not-support, PS: partial-support, IS: indirect-support, AQE: adaptive query execution, and Dynamic: dynamically changing operators and pipelines. Exe. is the abbreviation for Execution.
@@
Query performance prediction is a widely needed task in admission control [26, 15], resource management [15, 25, 27], and query optimization [4, 28, 39, 38], which mainly aims to estimate the latency of queries. The accuracy of QPP directly impacts all the above data management tasks and is crucial to the performance and stability of the DBMS. QPP can be regarded as a specific type of cost estimation problem that uses query latency to measure query cost. Before the widespread introduction of deep learning into the database domain, mainstream databases primarily employed conventional methods to predict query plan performance. These methods mostly relied on simple linear models, combining the cost coefficients of physical operators such as sequential scan and indexing with the input scale provided by cardinality estimators to derive query performance. Although such conventional methods are widely used in some advanced OLAP DBMSs such as SparkSQL [39] and Databricks [38], they ignore the inherent complexity of DBMS and can only achieve limited accuracy. For example, the single operator’s cost can be variable under different query pipeline contexts and different system workloads due to the cache changing and resource utilization caused by other operators, especially under the dynamic pipeline environment. Such complexity is even more significant in some OLAP DBMSs, which take the high degree of parallelism of query pipelines as the core design idea. Solving such complex problems is exactly what the learning cost estimator is good at.
@@
With the development of deep learning, numerous studies have applied deep learning models to the QPP field, aiming to directly measure the quality of query plans based on the time cost of query execution. Existing works primarily focus on feature representation and end-to-end prediction for tree-shaped query plans. For instance, in databases like PostgreSQL [33] where query plan trees serve as physical plans, many methods [4, 13, 11] have been demonstrated to achieve higher accuracy and prediction speed. In the field of concurrent query performance prediction, there have also been studies [12] treating concurrent queries’ plan trees as Directed Acyclic Graphs (DAGs) and treating resource competition and data sharing relationships between concurrent queries as edges that connect those sub-DAGs to capture resource competition relationships among a group of operators.
However, most existing learning-based methods are designed for DBMSs with tree-shaped query plans and static serial pipelines. They lack sufficient targeted design and awareness of the underlying query execution mechanisms in DBMSs, particularly for mechanisms commonly used in OLAP databases like vectorized operators, Directed Acyclic Graph (DAG) pipelines, parallel query execution engines, dynamic pipeline modifications (or adaptive query execution), as well as the synergistic effects of these mechanisms.
C.1. Difficulty in capturing the synergistic effects of complex execution mechanisms. An increasing number of database systems [20, 40, 31] employ SIMD instruction sets (such as SSE and AVX  [^1] https://clickhouse.com/blog/cpu-dispatch-in-clickhouse) to implement vectorized physical operators. These operators, due to their higher computational throughput, exert greater pressure on resources like CPU caches and memory I/O. When combined with mechanisms like parallel query execution, this makes it even more challenging for models to capture the key characteristics of query execution in an end-to-end manner, thereby limiting the accuracy of cost estimation.
@@
C.2 Dynamic pipeline extraction and expression. Some DBMSs [39, 38, 20] employ runtime optimization to alleviate issues such as inaccurate estimation of key information, such as cardinality, before querying, or changes in data distribution and system load during runtime. Such a mechanism will add extra operators during execution and change the operator algorithms, such as sort and join. In summary, there are significant differences between such high-performance OLAP DBMSs and other databases, and existing QPP methods are not designed for them. Such a mechanism makes it inaccurate to use the pipeline extracted before the query, and it is also difficult to express any dynamic pipeline extracted in a form that can be easily processed by learned QPP methods. It’s difficult to extract a low-level runtime dynamic pipeline that is detailed enough to support the QPP task. Also, even if such information can be extracted, it is still difficult to process with low cost by learned QPP methods, especially when the pipeline is in spatiotemporal graph form since the inference on spatiotemporal graph networks is expensive.
@@
C.3 Parallel execution and resource competition. Some DBMSs [39, 20] employ parallel pipeline executors, which introduce more complex resource competition within a single OLAP query. Taking ClickHouse  [^2] https://github.com/ClickHouse/ClickHouse/issues/34045 as a representative example, its parallel execution of ClickHouse consists of two parts, horizontal and vertical parallel. The horizontal parallel execution means that different parts of the data are processed in parallel by the same computation. ClickHouse achieves this by widely using SIMD instruction sets like SSE and AVX  [^3] https://clickhouse.com/blog/cpu-dispatch-in-clickhouse and creating multiple execution branches (threads) for operators like MergeSorting to speed up queries. Such operators’ costs are highly related to the system workload and much harder to predict within an end-to-end model. The vertical parallel execution means that different stages of computation are run in parallel. To fit such a design and speed up queries, different from the widely used volcano model, these DBMSs’ query plans are constructed as parallel pipelines [^4] https://presentations.clickhouse.com/meetup24/5.%20Clickhouse%20query%20execution%20pipeline%20changes/, resulting in significant resource competition among operators within a single query plan, including competition for CPU, I/O, and memory resources. Existing methods focus on common databases with tree-like query plans, how to capture the resource competition information, and calibrate operators’ costs still remain to be studied.
@@
In this paper, we propose a multi-stage QPP method, CONCERTO, to tackle the challenges above and give an accurate and efficient prediction, which decouples the operator cost prediction, cost calibration, and query performance prediction. The core idea of CONCERTO to achieve high accuracy for high-performance OLAP DBMSs is first to obtain detailed low-level physical pipeline information that includes dynamic pipeline changes, and then to use multi-stage models to predict the resource cost of SIMD operators separately, calibrate the resource competition, and summarize the overall pipeline latency.
@@
For challenge C.1, CONCERTO decouples operator cost prediction from query performance prediction. We establish lightweight cost models (a.k.a operator cost predictors (OCPs)) for each operator to predict the resource and time cost of each operator’s execution as Stage 1. These individual operator cost vectors are then aggregated into a graph for subsequent processing.
@@
For challenge C.2, CONCERTO includes a Runtime Tracker to extract training data and necessary dynamic information for prediction. To collect training data of the OCPs, CONCERTO uses a novel serial executor, and the tracker runs in full-collection mode that marks all data chunks by their addresses. For query-level training data of training and prediction, considering high-performance OLAP DBMSs’ dynamic parallel execution mechanism and inspired by ClickHouse’s probe phase [20], the tracker uses a probe execution mode which only launches a few chunks to obtain operator features and the execution paths, which usually contain more operators than the initial pipeline. For dynamic pipeline expression, CONCERTO uses the data-flow tree rather than the pipeline DAG as the backbone of the data structure. CONCERTO includes a Graph Constructor to build the resource competition graph with the backbone data-flow tree and corresponding operators’ cost vectors based on the execution paths of chunks collected by the Runtime Tracker. Therefore, the chunks’ execution information generated by the modified pipeline can be reserved in different execution paths of the data-flow tree without mixing up.
@@
For challenge C.3, we introduced resource competition into the pipeline and designed an attention-based resource competition intensity adaptive mechanism and then used Graph Attention Network(GAT) [21] to calibrate the cost as Stage 2. In particular, CONCERTO categorizes operators based on their resource usage and assigns them to different resource competition groups. We establish edges in the generated data-flow tree based on those resource competition groups to represent resource competition relationships among operators within the same group. The edges’ weights are self-adjusted by an attention mechanism that is coupled with the data flow structure. Next, we use a GAT model to calibrate operator costs. Finally, to gather those calibrated costs into the final QPP result, we develop a method to convert the graph back to a tree in a differentiable way and use a TCN to aggregate those costs from the bottom to the top and predict the query performance as Stage 3.
@@
    We designed a multi-stage query performance prediction approach that decouples operator cost estimation, resource contention calibration, and query performance prediction. This method provides more accurate query performance forecasting for database systems with complex underlying query mechanisms (such as high-performance OLAP), while also filling the research gap in learning-based cost estimation for database systems featuring dynamic parallel DAG pipelines.
  2. 2.
@@
    We design a low-level Runtime Tracker and implement it on ClickHouse to collect detailed and accurate data for the QPP model’s training and prediction, which includes full-collection mode, probe execution mode, Cost Logger, and a modified serial executor.
  3. 3.
@@
    Experimental results demonstrate that CONCERTO achieves higher estimation accuracy and comparable or lower time overhead than baseline methods.
II Related Works
II-A Conventional Methods
@@
The most common method is to build cost models for different operators before learning methods are introduced into the database system. Such methods employ the CPU cost, I/O cost, etc. of operators to estimate the query cost. Manegold et al. proposed a series of cost functions for different memory access patterns [1]. These functions are parameterized to fit different hardware features. The operators’ cost can be inferred from the combination of those cost functions. Feilong Liu et al. proposed a memory I/O cost model [2] to identify good evaluation strategies for complex query plans with multiple hash-based equip-joins over memory-resident data. Wentao et al. proposed a cost model [14] that represents the query latency as a function of the selectivities of operators in the query plan as well as the constants that describe the cost of CPU and I/O. They solve the coefficient of the cost model from the observed query information. Wu et al. [22] proposed an offline profiling method to set the coefficients of a cost model for specific hardware and software conditions.
@@
Such methods have two main problems, the first problem is that the parameters of the cost models are hard to adjust properly to maximize accuracy. Although there are some methods  [14, 37] proposed to adjust or solve the parameters, it’s not accurate enough for dynamic workload environments since the ground truth costs change with the hardware workload. The second problem is that some research  [3, 13] show that the cardinality estimation results are crucial to those cost models, which is another harder problem that usually has much higher error. One important reason is that, compared to learned models, conventional methods’ prediction results highly depend on the input scale and lack robustness.
II-B Learned Methods